OpenStudy (helpmeee123):

Rectangle dimensions (inches) Side a: side a +b : a/b: (a+b)/a: golden ratio? 5 8 5/3= 1.6 8/5=1.6 yes 3 4 6 10 14 22 12 16 15 24 What is the equation? And the answers? Please help

OpenStudy (helpmeee123):

Please....

jimthompson5910 (jim_thompson5910):

if a = 5 and a +b = 8 then a+b = 8 5+b = 8 b = 8-5 b = 3

jimthompson5910 (jim_thompson5910):

so that's how they got a/b = 5/3 = 1.6

Directrix (directrix):

Rectangle dimensions (inches) Side a: side a +b : b a/b: (a+b)/a: golden ratio? 5 8 3 5/3= 1.6 8/5=1.6 yes 3 4 1 6 10 4 14 22 8 12 16 4 15 24 9 Because you have to get the ratio of a/b, I added a b column to the chart. The b values I entered come from subtracting the given a value from (a + b).

Directrix (directrix):

@ Your task is to fill in the a/b and (a + b)/a columns. So do that now and post the results.

Directrix (directrix):

@helpmeee123

OpenStudy (helpmeee123):

Thankyou. :)

Directrix (directrix):

Fill in the chart and post the results. From that, we will figure out the equation.